Hello, this is a Micro Economic problems Could you please be kind enough and solve all of problems with the explanation in detail? Thank you and have a good one! 1. Suppose a firm has market power and faces a downward sloping demand curve for its product, and its marginal cost curve is upward sloping. If the firm reduces its price, then: A) consumer and producer surplus must increase. B) consumer surplus increases, producer surplus may increase or decrease. consumer surplus increases, producer surplus must decline. D) consumer and producer surplus must decline. 2. Suppose that the marginal cost of an additional ton of steel produced by a Japanese firm is the same whether the steel is set aside for domestic use or exported abroad. If the price elasticity of demand for steel is greater abroad than it is in Japan, which of the following will be correct? A) The Japanese firm will sell more steel abroad than they will sell in Japan. B) The Japanese firm will sell more steel in Japan than they will sell abroad. C) The Japanese firm will sell steel at a lower price abroad than they will charge domestic users. D) The Japanese firm will sell steel at a higher price abroad than they will charge domestic users. Your local grocery store offers a coupon that reduces the price of milk during the coming week. The regular retail price of milk in the store is $3.00 per gallon, and the coupon price is $2.00 per gallon for the next week. If the store maximizes profits and the price elasticity of demand for milk is -2 for coupon users, what is the price elasticity of demand for non-users? A) -0.67 B)-1.0 C)-1.5 D) We do not have enough information to answer the question. 3.
A firm has two customers with non-identical demands and a constant marginal cost of production. At any positive price, the consumer surplus values for the two customers are related as CS2CS1. What can we say about the optimal two-part tariff for the firm? A) The firm sets the price equal to MC and the optimal tariff is equal to CS2 B) The firm sets the price equal to MC and the optimal tariff is equal to CS1. C) The firm sets the price equal to MC and the optimal tariff is equal to zero. D) The optimal price is greater than MC and the optimal tariff is equal to CS1. 4. 5. One Guy's Pizza advertising expenditures are $1,200 and sales are $30,000. When the advertising expenditure increases to $1,400, pizza sales increase to $32,000. The arc advertising elasticity of demand is approximately A) 0 B) 0.1 C) 0.4 D) 2.5 I will give you thumb up!
Answer #1

Ans:1 C) consumer surplus increases, producer's surplus must decline.

Reason: the condition mentioned in the question is called Monopoly market structure. A monopolist charges a higher price to increase its profit and producer's surplus. Any decrease in price will increase consumer's surplus but will have an ill effect on the producer's surplus.

Ans:2 C) the Japanese firm will sell steel at a lower price abroad than they will charge on domestic consumers.

Reason: with increased elasticity in abroad, the percentage change in price will be less than the percentage change in quantity demanded. this means, if the firm slightly increases the price, this decreases the higher quantity of demand in abroad than in the domestic market.

Ans:3 D) we do not have enough information to answer this question

Ans:5 C) 0.4

Reason: Change in quantity demanded= (30,000-32000) / 30,000 = -0.06

Change in advertising expenditure= (1,200- 1,400)/ 1,200 = -0.17

Arc elasticity= change in quantity/ change in price = -0.06/-0.17= 0.35 or approx 0.4